### v3

# Can Large Language Models Master Complex Card Games?

Abstract
Complex games have long been an important benchmark for testing the progress of artificial intelligence algorithms. AlphaGo, AlphaZero, and MuZero have defeated top human players in Go and Chess, garnering widespread societal attention towards artificial intelligence. Concurrently, large language models (LLMs) have exhibited remarkable capabilities across various tasks, raising the question of whether LLMs can achieve similar success in complex games. In this paper, we explore the potential of LLMs in mastering complex card games. We systematically assess the learning capabilities of LLMs across eight diverse card games, evaluating the impact of fine-tuning on high-quality gameplay data, and examining the models’ ability to retain general capabilities while mastering these games. Our findings indicate that: (1) LLMs can approach the performance of strong game AIs through supervised fine-tuning on high-quality data, (2) LLMs can achieve a certain level of proficiency in multiple complex card games simultaneously, with performance augmentation for games with similar rules and conflicts for dissimilar ones, and (3) LLMs experience a decline in general capabilities when mastering complex games, but this decline can be mitigated by integrating a certain amount of general instruction data. The evaluation results demonstrate strong learning ability and versatility of LLMs. The code is available at https://github.com/THUDM/LLM4CardGame
11footnotetext: Corresponding author.
1 Introduction
A long-term goal of artificial intelligence is to achieve superhuman performance in highly challenging domains [1, 2, 3]. Games, particularly complex ones such as Chess and Go, have become the best testing grounds for artificial intelligence algorithms [4, 5, 6, 7]. In recent years, artificial intelligence algorithms have made significant breakthroughs in the realm of games. AlphaGo is the first to defeat human professional players in Go by using supervised learning from expert human data and reinforcement learning [4]. Following this, a general reinforcement learning algorithm, AlphaZero, achieves superhuman performance in three challenging games: Chess, Shogi (Japanese chess), and Go [6]. MuZero even achieves performance equivalent to AlphaZero without needing to know the rules of the game [7].
Recently, large language models (LLMs) [8, 9, 10, 11] have achieved remarkable performance, even surpassing human levels, across a wide range of tasks including general knowledge question answering [12, 13], mathematics [14, 15], coding [16, 17], and agent [18, 19]. This naturally raises the question: can language models achieve superhuman performance in complex games, or at least reach the same level as the best reinforcement learning algorithms? In this paper, we primarily focus on card games.
Currently, there is a vast amount of research evaluating LLMs across various tasks. Among these studies, some research also evaluates the decision-making capability of LLMs in both board games and card games, such as Texas Hold’em [20, 21], Blackjack [22], and Guandan [23, 24]. However, these studies still have some limitations. First, many evaluation studies assess LLMs through prompting without involving fine-tuning [25, 26]. These prompt-based evaluation studies can only assess whether LLMs are capable of applying their existing knowledge to new environments. But, they do not evaluate the learning ability of LLMs. Second, some evaluation studies include assessments of LLMs after fine-tuning and demonstrate that fine-tuning improves the performance of LLMs in new environments. But, the tasks evaluated in these studies lack sufficient complexity, making them inadequate to comprehensively assess the learning capabilities of LLMs.
Figure 1: Complexity of the games. InfoSet Number: the number of the information sets; Avg. InfoSet Size: the average number of states in a single information set.
As mentioned at the beginning, complex games are often used to explore the upper bound of artificial intelligence algorithms’ learning capabilities. Therefore, in this paper, we investigate whether language models can master complex card games. To address the shortcomings of previous work, we systematically evaluate the performance of language models on eight carefully selected card games. First, most games exhibit a high level of complexity as shown in Figure  1. The high complexity of these games presents a greater challenge to the learning abilities of large models. Evaluations across the eight games provide a more comprehensive understanding of LLMs. Second, we evaluate the performance ceiling that LLMs can achieve by fine-tuning the model on high-quality gameplay interaction data. Compared to prompt-based methods, fine-tuning methods focus on evaluating the learning ability of language models. We generate high-quality gameplay interaction data using strong game AIs or directly utilize publicly available high-quality interaction data. Specifically, we focus on the following three research questions:
1. Can LLMs master complex card games? And how much data is required to master these games?
2. Can LLMs simultaneously master multiple games? Do different games mutually enhance each other or do conflicts arise between them?
3. Can LLMs maintain their general capabilities while mastering complex games?
To answer these questions, we first fine-tune language models on each of the eight games separately to evaluate the extent to which the models can master individual games. Next, we fine-tune the models on a mixture of all the game data to assess their ability to master all the games simultaneously. Finally, we evaluate whether the models’ general capabilities decline using MMLU-Pro [12], Math-500 [27], and HumanEval [28] benchmarks for knowledge question answering, math, and coding skills. Additionally, we analyze the performance variations of language models with different parameter sizes and types (Qwen2.5 [29], Llama3.1 [30], and GLM4 [9]). In summary, the contributions of this work are:
  • We are the first to propose a comprehensive evaluation of the learning capabilities of LLMs across multiple high-complexity games, which present greater challenges to the learning abilities of LLMs.
  • We obtain a large amount of high-quality data for LLMs to learn by utilizing strong game AIs and game prompt templates, avoiding the problem of high computational resource consumption when LLMs learn from scratch in the environment.
  • We thoroughly assess the learning ability of the models through experiments in three aspects. The evaluation results demonstrate strong learning ability and versatility of LLMs, as they can simultaneously master multiple complex games without altering the model structure.
2 Related Works
The integration of LLM-based agents into card and board game has garnered significant attention, focusing on enhancing strategic reasoning and adaptability [31, 32, 33]. Previous research has mainly focused on improving LLM-based agents in card games with incomplete information and adversarial settings, such as Texas Hold’em [20, 21, 22, 34], Blackjack [22], and Guandan [23, 24]. For example, Yim et al. [23] introduced a Theory of Mind (ToM) planning technique for Guandan, enabling LLM agents to adapt strategies based on game rules, state, and history. Their approach also integrates an external tool to manage the game’s large action space. While effective in a single game, the generalizability of such prompt-based strategies across multiple games remains an open question. Guo et al. [21] proposed Suspicion-Agent, a prompt-based approach that leverages GPT-4’s reasoning and high-order ToM capabilities to adapt strategies in imperfect information card games. It enables GPT-4 to play against different opponents using only game rules and observations as input. Their experiments demonstrate its effectiveness across multiple games. However, as a prompt-based method, it relies solely on the model’s inherent knowledge, which limits its overall performance.
3 Method
From AlphaGo, to AlphaZero, and then to MuZero, we can see that these methods have achieved significant breakthroughs in complex games by continuously exploring the environment and leveraging successful experiences. This paper aims to explore whether LLMs can master complex games similarly. Considering the time and resource consumption involved in the exploration process, we utilize existing strong game AIs to generate high-quality trajectory data. This study investigates whether LLMs can master complex games by learning from this high-quality trajectory data. Next, we introduce the selected games and the process of generating training data for each game.
3.1 Games
For game selection, we primarily consider the popularity, complexity, and availability of high-quality models or data. Based on these three aspects, we select eight games: DouDizhu, Guandan, Riichi Mahjong, Uno, Gin Rummy, Leduc Hold’em, Limit Texas Hold’em, and No-Limit Texas Hold’em.
DouDiZhu. DouDizhu [^1] https://en.wikipedia.org/wiki/Dou_dizhu (a.k.a. Fighting the Landlord) is the most popular card game in China. The game is played by three players with a standard 54-card deck. There are two roles in the game: a landlord and two farmers. Some studies have explored building strong DouDizhu AIs using techniques such as handcrafted heuristic rules, reinforcement learning based methods, and search algorithms [35, 36, 37]. Among these methods, DouZero [37] is a simple and effective approach that requires no human knowledge or state/action abstraction. It is currently the strongest publicly accessible DouDizhu AI.
GuanDan. Guandan [^2] https://en.wikipedia.org/wiki/Guandan is another popular card game in China. The game requires four players, with the two players sitting opposite each other forming a team. Its gameplay is similar to Dou Dizhu. However, Guandan is more complex in comparison. There is relatively less research on Guandan AI [38, 39]. Among the existing work, DanZero [38], which employs a neural network-enhanced Monte-Carlo method, has outperformed other algorithms. Therefore, we choose DanZero as our teacher model.
Mahjong. Mahjong [^3] https://en.wikipedia.org/wiki/Mahjong is a widely popular multiplayer tile-based game across the world. Mahjong has many variants, and this paper focuses on Riichi Mahjong (a.k.a. Japanese Mahjong). Suphx is the first Mahjong AI to defeat most top human players [40]. Then LuckJ developed by Tencent reached 10 Dan on Tenhou [^4] https://tenhou.net/ and surpasses all human players and other AIs. However, the model weights for both of these AIs have not been made publicly available. Nevertheless, Tenhou provides gameplay data from expert players.
Uno. Uno [^5] https://en.wikipedia.org/wiki/Uno_(card_game) is a proprietary American shedding-type card game. The game is played with a specially designed 108-card deck. There are 2 players in the game. Each player starts with seven cards dealt face down. Players take turns matching the card in the Discard Pile by number, color, or symbol/action. The objective is to be the first player to get rid of all the cards in hand.
Gin Rummy. Gin Rummy [^6] https://en.wikipedia.org/wiki/Gin_rummy is a two-player card game. The game is played by two players using a standard 52-card deck. The dealer deals 11 cards to the opponent and 10 cards to himself. During each turn, you can pick up the discard or draw from the face-down stock, then discard a card. Players try to form melds of 3 or more cards of the same rank or 3 or more cards of the same suit in sequence.
Game # Players # Teams Teacher model/Data Opponent # Games
Avg. Steps
per Game
Avg. Steps
per Player
per Game
Total Steps
Avg. Legal
Actions
per Step
Training data
DouDizhu 3 2 DouZero Rule model 200k 37.31 12.44 2,950k 10.06 1,000k
GuanDan 4 2 DanZero Rule model 6k 311.25 155.63 1,220k 48.67 1,000k
Riichi Mahjong 4 4 Data from experts Mortal 7k 656.92 164.23 1,090k 8.79 1,000k
Uno 2 2 Rule model Random 50k 42.33 21.16 410k 3.14 400k
Gin Rummy 2 2 Rule model Random 50k 52.14 26.07 1,280k 6.22 400k
Leduc Hold’em 2 2 DQN model Random 400k 3.61 1.81 580k 2.86 400k
Limit Texas Hold’em 2 2 DQN model Random 200k 5.01 2.50 450k 2.96 400k
No-limit Texas Hold’em 2 2 DQN model Random 400k 3.78 1.89 700k 4.31 400k
Table 1: Data generation information of games.
Leduc Hold’em. Leduc Hold’em, introduced in  Southey et al. [41], is a simplified variant of Limit Texas Hold’em. This version uses a deck comprising only six cards, with two pairs each of King, Queen, and Jack. The game involves two players, spans two rounds, and has a maximum of two bets per round, with the raise amounts fixed at 2 in the first round and 4 in the second round.
Limit Texas Hold’em. Limit Texas Hold’em [^7] https://en.wikipedia.org/wiki/Texas_hold_%27em is a well-known betting game with 52-card deck. Each player receives two face-down cards, known as hole cards. Subsequently, five community cards are dealt in three stages: the flop, the turn, and the river. Players aim to form the best possible five-card hand using any combination of their hole cards and the community cards.
No-limit Texas Hold’em. No-limit Texas Hold’em follows similar rules to Limit Texas Hold’em but with key differences in betting. No-limit Texas Hold’em allows players to raise by at least the amount of the previous raise in the same round and up to the entirety of their remaining stack. Additionally, there is no limit on the number of raises in No-limit Texas Hold’em.
3.2 Data Preparation
Trajectory Generation. We generate gameplay interaction data by having the teacher model compete against opponents. The teacher model and opponent information for each game are shown in Table 1. For DouDizhu, we use DouZero [37] as the teacher model and a rule-based model [42] as the opponent model. For GuanDan, we use DanZero [38] as the teacher model and a rule-based [38] model as the opponent model. For Riichi Mahjong, we download the match data of human professional players from the Tenhou [^8] https://tenhou.net/ platform for the year 2020. The opponent model is Mortal [^9] https://github.com/Equim-chan/Mortal, a strong Mahjong AI, which is used only during evaluation. For Uno and Gin Rummy, we use rule model from  Zha et al. [42] as the teacher model and use random as the opponent. For Leduc Hold’em, Limit Texas Hold’em, No-limit Texas Hold’em, we train DQN model as the teacher model with RLCard framework [^10] https://github.com/datamllab/rlcard.
Based on the complexity of different games, we play each game a varying number of times. The important information of the generated data is shown in Table 1. From the table, it can be seen that the average number of steps in the games Doudizhu, Guandan, and Mahjong is significantly higher than the other games. Particularly, Guandan and Mahjong have longer steps because each game consists of multiple rounds. For example, in Guandan, the game progresses from 2 to Ace.
Trajectory Filtering. In this paper, we consider each observation-action pair of one step as a sample. We filter the generated data based on two criteria to obtain high-quality data. First, we only retain the observation and action data of the winning player. Additionally, we consider each observation-action pair from the player as a single data instance. Second, for all eight games, the environment provides the legal action options per action, and we only retain data samples where the number of legal action options is greater than one. The amount of data obtained after filtering is presented in Table 1. It can also be seen from the table that the number of legal candidate actions per sample for Doudizhu, Guandan, and Mahjong exceeds that of the other five games, making these three games relatively more complex.
Supervised Fine-Tuning Data Generation. To perform instruction fine-tuning on the model, we design prompts for each game to convert observation-action pairs into instructions and corresponding outputs. The instruction primarily consists of three parts: game introduction, state data, and output format instructions. The game introduction includes the game rules and the player’s objectives. The state data comprises information such as the player’s hand, community cards, the sequence of historical actions, and legal actions. The output format specifies that the model should output actions in JSON format. Complete instructions for each game can be found in Appendix  A.3.
4 Experiments
4.1 Experiment Setup
Data. Through the data synthesis process described in Section 3.2, we obtain the training data for each game. For DouDizhu, GuanDan, and Riichi Mahjong, we sample 1,000k instances as training data. For Uno, Gin Rummy, Leduc Hold’em, Limit Texas Hold’em, and No-limit Texas Hold’em, we sample 400k instances as training data.
Model. We fine-tune three different types of language models—Qwen2.5-7B-Instruct [29], Llama3.1-8B-Instruct [30], and GLM4-9B-Chat [9]—to analyze the impact of model type on performance. Additionally, we fine-tune Qwen2.5 models with different parameter sizes, ranging from 0.5B to 14B parameters, to evaluate the effect of model size on performance. We fine-tune all models with LLaMA-Factory Framework [43] and use LoRA fine-tuning [44]. The LoRA rank and LoRA alpha are set to 8 and 16, respectively. We fine-tune all models with 1 epoch. We apply a peak of 1e-4 learning rate with a cosine scheduler. The batch size is 128. We conduct experiments on a server with 8 H100 GPUs.
Metric. For DouDizhu, we use the win rate. For GuanDan, we use the round win rate. For Riichi Mahjong, Uno and Gin Rummy, Leduc Hold’em, Limit Texas Hold’em, No-limit Texas Hold’em, we use reward score. The reward score in Mahjong is determined based on the average rank over multiple games, with the rewards for ranks 1, 2, 3, and 4 being 3, 2, 1, and 0, respectively. The reward scores for the other five games can be found in the RLCard framework [^11] https://rlcard.org/. We evaluate the LLM by having it play multiple games against opponents. The number of games for the eight games are 1000, 20, 50, 500, 100, 1000, 1000, and 1000, respectively. For DouDizhu, following DouZero [37], we have the LLM play 500 times as the Landlord and 500 times as the Farmers, then calculate the average win rate for both roles.
4.2 RQ1. Can LLMs master complex card games? And how much data is required for them to master these games?
Experimental Design. We fine-tune the language model separately on each game’s data and then evaluate its performance on the respective game. The training data information for each game is presented in Table 1. For each game, we train on all training data for one epoch, saving a checkpoint every 400 steps. This allows us to analyze the model’s performance changes with different amounts of training data. Additionally, we fine-tune three different types of LLMs and five different sizes of LLMs to explore the impact of model type and size on performance.
(a) DouDizhu
(b) GuanDan
(c) Riichi Mahjong
Figure 2: Performance of different training data.
Results and Analysis. The results of DouDizhu, GuanDan, and Mahjong are shown in Figure 2(a)-Figure 2(c). As shown in the figure, with the increasing amount of training data, the performance of the LLM in Doudizhu and Guandan gradually approaches that of the teacher model. In Mahjong, even though there is no available teacher model, the LLM has achieved performance comparable to that of a very strong Mahjong AI. Figure 1 and Table 1 have already shown that these three games have high complexity (long average decision steps and a large number of valid actions per step). These results indicate that, given sufficient high-quality data, LLMs can master complex card games. As training progresses, the model acquires more and more game strategic knowledge, leading to a continuous improvement in win rate. It is worth noting that DouZero actually consists of three models, with one model trained for the Landlord and two models for the two Farmers. In contrast, LLMs can play all three roles with a single model, further demonstrating their powerful learning capability and versatility.
(a) DouDizhu-Landlord
(b) DouDizhu-Farmer
Figure 3: Performance of different roles.
Figure 4: Performance of models with different sizes on DouDizhu.
Different Games and Model Types. As shown in Figure 2(b) and Figure 2(c), in Guandan and Mahjong, there is no significant difference in performance among the three models, indicating that the learning capabilities of the three models are comparable. However, in DouDizhu, the performance of GLM is significantly worse than Qwen and Llama. To analyze the performance differences of different models in DouDizhu, we further plot the win rates of the models when playing different roles, as shown in Figure 3(a) and Figure 3(b). As observed from Figure 3(a) and Figure 3(b), GLM performs better than Qwen and Llama in DouDizhu-Landlord, while performing significantly worse than the two models in DouDizhu-Farmer. This suggests that GLM did not effectively balance the learning between the two roles and focused more on the landlord role, leading to weaker performance for the farmer role.
However, comparing Figure 3(a) and Figure 3(b), we discovered another strange phenomenon: for DouDizhu, why is there such a difference between Landlord and Farmer performance? And why are all of the models so much lower than ceiling performance for the Farmer role? Upon analyzing the training data, we suspect that this is caused by the filtering rules applied to the game data. In DouDizhu, there is one landlord and two farmers, with each game resulting in either a victory for the landlord or for the farmers. During filtering, we retained only the data for the winning side. When the farmers win, the data for both farmers is retained. However, in many cases, the victory may have been primarily driven by the actions of one farmer, while the data for the other farmer may be of lower quality. Consequently, the training data includes some low-quality data for the farmer role, leading to performance for the farmer role that falls significantly below its ceiling performance.
Model Landlord Farmer Average
Qwen2.5-7B-Instruct 0.828 0.784 0.806
Qwen2.5-14B-Instruct 0.858 0.570 0.714
Table 2: Win rate of 7B and 14B models.
Different Model Sizes. We train and evaluate five different sizes of Qwen2.5 models on Doudizhu and the results are shown in Figure 4. From 0.5B to 7B, the performance of the models gradually improves as the number of model parameters increases, indicating a positive correlation between model size and performance. However, we notice that despite the 14B model having the largest number of parameters, its performance is the worst.
In order to analysis this, we further plot the win rates of the models when playing different roles in Table 2. From the table, 14B model performs better as the landlord (approaching the performance of the teacher model) but significantly worse as the farmer. This results in the average win rate of the 14B model being lower than that of the 7B model. This is similar to why the GLM model performs worse on DouDizhu compared to Qwen and Llama.
4.3 RQ2. Can LLMs simultaneously master multiple games? Do different games mutually enhance each other or do conflicts arise between them?
Experimental Design. Based on the above experimental results, we roughly determine the amount of data required for each game to converge. We then sample data from the training datasets of each game according to this amount and merge them to obtain a mixed training set that includes data from all games. Specifically, the combined dataset contains 3.1 million data points, with the quantities of the eight games being: 700k, 950k, 650k, 200k, 50k, 250k, 200k, and 100k, respectively. We empirically determine the quantity of instances for each game based on game complexity and the results of Experiment 4.2. For example, games with higher complexity necessitate a larger volume of training data. We fine-tune the language model on this mixed training set to evaluate whether it can simultaneously master multiple games. We compare the fine-tuned models with the API-based models and the base models.
Model DouDizhu GuanDan Riichi Uno Gin Rummy Leduc Limit Nolimit
API-based models
GPT-4o-mini 0.195 0.019 0.15 0.128 -0.176 0.30 0.45 2.47
GPT-4o 0.180 0.019 0.25 0.072 0.405 0.84 0.60 2.73
GLM-4-air 0.330 0.000 0.10 -0.068 -0.415 -0.12 1.13 -0.89
GLM-4-plus 0.345 0.019 0.00 0.020 -0.344 0.80 0.63 3.21
DeepSeek-V3 0.320 0.000 0.15 0.128 0.147 0.77 0.22 0.18
DeepSeek-R1 0.185 0.020 0.05 0.148 0.228 0.88 0.24 1.88
Base models
Qwen2.5-7B-Instruct 0.087 0.000 0.04 0.032 -0.530 0.63 1.05 1.25
Llama3.1-8B-Instruct 0.155 0.000 0.08 0.120 -0.463 0.62 -0.04 -2.10
GLM4-9B-Chat 0.131 0.000 0.08 0.000 -0.362 0.52 0.85 -0.44
Fine-tuned models
Qwen2.5-7B-Instruct-mix 0.852 0.634 1.08 0.108 0.177 1.24 2.66 4.86
Llama3.1-8B-Instruct-mix 0.870 0.661 1.38 0.164 0.186 1.24 2.77 6.02
GLM4-9B-Chat-mix 0.882 0.698 1.31 0.252 0.191 1.24 2.89 5.77
Table 3: Performance of different models on all games. Bold font indicates the maximum value in each category, and underline indicates the second-highest value. Mix refers to models fine-tuned on the mixed training set composed of data from all games.
Results and Analysis. The results are shown in Table 3. All API-based models score relatively low on the two most complex games, GuanDan and Riichi, while their scores are relatively higher on the other six games. DeepSeek-R1 performed the best among all API-based models, achieving the highest scores in three games. This demonstrates the effectiveness of the reasoning mode. We observe that models of the same type with larger parameter versions or reasoning versions perform better than those with smaller parameters or non-reasoning versions. For example, DeepSeek-R1 shows improvements in most games compared to DeepSeek-V3. GLM and DeepSeek’s models score higher in DouDizhu, likely because this game is quite popular in China. Compared to the API-based models, the three base models perform worse in most of the games. Compared to the API-based models and base models, the fine-tuned model achieve the best performance in most of the games. Notably, in the complex games of Doudizhu, Guandan, and Riichi, their performance improve significantly. These results indicate that, after being trained on multiple high-quality game datasets, LLMs can simultaneously master multiple complex games.
Model/Game DouDizhu GuanDan Riichi Uno Gin Rummy Leduc Limit Nolimit
DouDizhu 0.806 0.010 0.08 0.032 -0.528 0.637 1.16 2.54
[…]
Riichi 0.022 0.010 1.44 0.000 -0.233 0.637 0.91 -0.87
Uno 0.101 0.000 0.06 0.220 0.028 0.637 1.14 1.45
[…]
Nolimit 0.118 0.000 0.10 -0.056 -0.432 1.012 2.12 7.75
Mix 0.852 0.634 1.08 0.108 0.177 1.244 2.66 4.86
Table 4: Influence between different games using Qwen model. Each row represents the performance of a model trained on one specific game across all games. Mix refers to models fine-tuned on the mixed training set composed of data from all games. Bold indicates the maximum value, and underline indicates the second-highest value, both excluding the mix model.
Model/Game DouDizhu GuanDan Riichi Uno Gin Rummy Leduc Limit Nolimit
DouDizhu 0.824 0.000 0.13 0.008 -0.496 0.637 1.14 3.21
[…]
Riichi 0.142 0.000 1.42 0.060 -0.242 0.757 0.95 -1.07
[…]
Nolimit 0.170 0.000 0.04 0.056 -0.198 1.000 2.06 4.92
Mix 0.870 0.661 1.38 0.164 0.186 1.244 2.77 6.02
Table 5: Influence between different games using Llama model.
Model/Game DouDizhu GuanDan Riichi Uno Gin Rummy Leduc Limit Nolimit
[…]
Uno 0.111 0.000 0.06 0.176 -0.302 0.637 1.14 2.20
Gin Rummy 0.075 0.000 0.06 0.016 0.196 0.637 1.12 2.88
[…]
Nolimit 0.114 0.000 0.06 0.004 -0.499 0.648 1.53 6.24
Mix 0.882 0.698 1.31 0.252 0.191 1.244 2.89 5.77
Table 6: Influence between different games using GLM model.
Influence Between Different Games. To explore the mutual influence between different games, we evaluate the model fine-tuned on one game across the other seven games. The results are shown in Table 4,5,6. Compared to models trained on other games, the model trained on GuanDan also performs well on DouDizhu. This indicates that GuanDan has a positive influence on DouDizhu. Additionally, we can see that there are also positive influences among the three games, Leduc Hold’em, Limit Texas Hold’em, and No-limit Texas Hold’em. We claim that if the rules of two games are more similar, there tends to be greater knowledge transfer between them, for example, DouDizhu and GuanDan, which have similar rules. Similarly, the three poker games, Leduc Hold’em, Limit Texas Hold’em, and No-limit Texas Hold’em, exhibit more knowledge transfer due to their similar game rules. Compared to DouDizhu and GuanDan, the rules of these three games are more similar, leading to a more significant transfer effect. Therefore, game rules primarily dictate knowledge transfer between different games.
Figure 5: Comparison between models fine-tuned on a single game and models fine-tuned on all games.
We also compare the models fine-tuned on a single game with those fine-tuned on all games. The comparison results for Doudizhu and Guandan are shown in Figure 5. Because the card-playing rules of Doudizhu and Guandan are very similar, the performance of the mixed fine-tuned models improves further on both games compared to the models fine-tuned on each game individually. This indicates that Doudizhu and Guandan can mutually enhance each other’s performance. However, we also observe that the performance of the mixed fine-tuned models declined on the other six games compared to the individually fine-tuned models. This suggests that there is a conflict between Doudizhu and Guandan and the other six games.
Figure 6: Performance of models with different sizes on DouDizhu and GuanDan.
Different Model Sizes. We train mixed models of different sizes on Qwen2.5. Figure 6 shows the performance of these models on Doudizhu and Guandan. The performance improves as the number of model parameters increases.
4.4 RQ3. Can LLMs maintain their general capabilities while mastering complex games?
Experimental Design. To test whether the models lose their general capabilities after mastering the games, we evaluate the models’ performance of general knowledge question answering, mathematics, and coding before and after fine-tuning, using MMLU-Pro [12], Math-500 [27], and HumanEval [28] benchmarks. If general capabilities decline after fine-tuning on games, can further fine-tuning on knowledge, mathematics, and coding data help restore these general capabilities to some extent?
Figure 7: Evaluation results of different models on general benchmarks.
Results and Analysis of General Benchmarks. The results are shown in Figure 7. The mixed models fine-tuned on all games show significant declines in their abilities in knowledge-based question answering, mathematics, and coding. We then further fine-tune the game model on a mixed dataset composed of knowledge data, mathematics data, coding data, and game data. The proportions of these four types of data were 20k, 20k, 20k, and 8k, respectively. The 8k game data consists of 8 games, with 1k data points for each game. The quantities and proportions are chosen by referring to previous work on general knowledge recovery [45]. The knowledge data, mathematics data, and coding data are taken from part of Tulu-3’s post-training data [46], as this model has made all its post-training data open source. The evaluation results of the model fine-tuned with the general data are shown in Figure 7. As shown in the table, by further fine-tuning on specific types of data, the model can restore its ability in specific areas to some extent, as demonstrated in the paper with knowledge-based question answering, mathematics, and coding capabilities.
Model DouDizhu GuanDan Riichi Uno Gin Rummy Leduc Limit Nolimit
Llama3.1-8B-Instruct-mix 0.870 0.661 1.38 0.164 0.186 1.24 2.77 6.02
[…]
GLM4-9B-Chat-mix 0.882 0.698 1.31 0.252 0.191 1.24 2.89 5.77
[…]
Table 7: Performance of fine-tuned models on games. Mix refers to models fine-tuned on the mixed training set composed of data from all games. General refers to models fine-tuned on the mixed training set composed of the knowledge data, mathematics data, and coding data.
Results and Analysis of Games. To evaluate the impact of general data fine-tuning on game performance, we provid the performance of the models on all games before and after fine-tuning with general data in Table 7. From the table, it can be seen that the model’s performance on games has remained mostly unchanged (slight improvements or stability in 5 games, and slight decreases in 3 games), indicating that the model regains a certain level of general capability while maintaining its gaming ability.
Different Model Types. In the three non-gaming benchmarks (MMLU-Pro, Math-500, and HumanEval), after fine-tuning on game data, GLM exhibites greater performance degradation compared to LLaMA. Furthermore, after fine-tuning on general-purpose data, GLM showes a lower degree of recovery on all three benchmarks relative to LLaMA, particularly on MMLU-Pro. This indicates that LLaMA is better than GLM at maintaining general capabilities, especially in retaining general knowledge. This may be related to the differences in training data and training methods used by the two models.
4.5 Discussion on the advantages of LLMs compared to specialized game AI
Comparison of computation and data. We want to compare the amount of computation and data required for fine-tuning versus training a game AI system from scratch. However, due to the insufficient information disclosed about these game AI systems and differences in hardware environments, conducting a fair comparison is infeasible. Nevertheless, we have tried to list some comparative information in Appendix A.1. This information does not directly demonstrate the advantages of LLMs in terms of computation and data requirements during training.
However, the key aspect we aim to highlight is that the greatest advantage of LLMs lies in their nature as general-purpose learners. To achieve strong performance in games, both traditional reinforcement learning approaches and LLMs require the selection of appropriate game features. In this regard, both approaches are similar. Nevertheless, traditional reinforcement learning methods require the design of network architectures that are tailored to the chosen features. Different games employ different features, meaning that each game necessitates a specially designed network architecture—a step that is often the most labor-intensive. LLMs, on the other hand, eliminate the need for this step of network structure design and can be applied across all games without modification. We only need to design templates for each game. This flexibility is the foremost advantage of LLMs. For example, DouZero, DanZero, and Mortal have network architectures specifically designed for individual games. DouZero even requires separate designs for the two roles in the game. In contrast, LLMs can perform well in three different games using the same architecture. Thus, we argue that the general learning ability of LLMs represents their most significant advantage.
5 Conclusion
In this paper, we explore the potential of large language models (LLMs) to master complex card games, evaluating their performance through fine-tuning on carefully selected high-quality gameplay interaction data. We explore three key research questions concerning LLMs’ ability to master complex card games, their capacity to learn multiple games simultaneously, and the impact of game mastery on their general capabilities. Our study reveals that LLMs have the potential to achieve strong performance in complex card games, while also handling multiple games at once and retaining significant portions of their general capabilities.
Acknowledgements
This research is supported by the National Natural Science Foundation of China (Grant No. 62406163).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We present our contributions in the Abstract and Introduction 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations in B.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: Our work is primarily an empirical analysis, and does not introduce theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have included all necessary information for reproducing the paper in the Experiment Setup 4.1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have provided the GitHub link to the code in the Abstract.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We have included all necessary information for reproducing the paper in the Experiment Setup 4.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: Error bars are not reported because it would be too computationally expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We listed the computational resources used in the Experiment Setup 4.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss societal impacts in Broader Impact C.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work evaluates the learning capabilities of large models through games and poses no risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: In our code repository, we referenced the creators’ code repository and used it appropriately according to its license.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We have provided detailed usage instructions in the code repository.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work does not involve crowdsourcing experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We detailed how to evaluate LLMs in the Experimental section 4.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Appendix
A.1 Comparison of computation and data
Among the three strong game AI models, Mortal, a Mahjong AI, does not have corresponding published papers, nor does its Git repository specify the computation and data required for training. Both DouZero and DanZero have published papers. Below is a comparison of our fine-tuned models with these two models in terms of hardware environment, training time, and data volume:
DouZero: According to their paper, DouZero was trained on a single server with 2 Intel(R) Xeon(R) Silver 4214R CPUs and 4 1080 Ti GPUs for 30 days; the data volume was not specified.
LLM-Dou: For the model in Figure 2(a) of our paper, we fine-tuned using a single server equipped with 2 Intel(R) Xeon(R) Platinum 8476C CPUs and 8 H800 GPUs on a dataset with 1 million samples. The fine-tuning times for the three models are as follows: Qwen2.5-7B: 11 hours; Llama3.1-8B: 12 hours; GLM4-9B: 14 hours.
DanZero: According to their paper, DanZero was trained on a server with 4 Intel(R) Xeon(R) Gold 6252 CPUs and a GeForce RTX 3070 GPU for 30 days; the data volume was not specified.
LLM-Dan: For the model in Figure 2(b) of our paper, we fine-tuned using a single server equipped with 2 Intel(R) Xeon(R) Platinum 8476C CPUs and 8 H800 GPUs on a dataset with 1 million samples. The fine-tuning times for the three models are as follows: Qwen2.5-7B: 21 hours; Llama3.1-8B: 25 hours; GLM4-9B: 29 hours.
A.2 Evaluation on more general benchmarks
We provide the results of the models on four other common benchmarks (GQPA-Diamond, AIME2024, LiveCodeBench, IFEval) before and after fine-tuning on general mixed data in Table 8. From the table, it can be seen that if the general mixed data does not include a specific type of data, the model’s corresponding capability will not be restored (after fine-tuning on the general mixed data, the performance of both models declined on the instruction-following benchmark).
Model MMLU-Pro Math-500 HumanEval GQPA-Diamond AIME2024 LiveCodeBench IFEval
Average
First_Three
Average
ALL
Llama-3.1-8B-Instruct 47.95 46.60 70.73 21.21 6.67 20.25 74.68 55.09 41.16
Llama-3.1-8B-Instruct-mix 44.74 35.20 60.98 26.77 6.67 17.75 74.31 46.97 38.06
Llama-3.1-8B-Instruct-mix-general 45.18 47.20 65.24 27.27 10.00 13.50 68.95 52.54 39.62
GLM-4-9B-Chat 48.04 52.40 76.22 26.26 0.00 18.00 69.13 58.89 41.44
GLM-4-9B-Chat-mix 35.12 37.20 39.63 26.26 0.00 13.75 63.40 37.32 30.77
GLM-4-9B-Chat-mix-general 31.75 41.00 68.90 20.20 0.00 16.25 56.01 47.22 33.44
Table 8: Evaluation results of different models on more general benchmarks.
A.3 Prompt Template
Figure 1: Prompt Template of DouDizhu Figure 2: Prompt Template of GuanDan Figure 3: Prompt Template of Riichi Mahjong Figure 4: Prompt Template of Uno Figure 5: Prompt Template of Gin Rummy Figure 6: Prompt Template of Leduc Hold’em Figure 7: Prompt Template of Limit Texas Hold’em Figure 8: Prompt Template of No-limit Texas Hold’em
Appendix B Limitations
Although the language models achieved performance close to that of strong game AIs, we found that the inference time of LLMs in games is relatively longer compared to these AIs. This is because these game AIs often have a smaller number of parameters, while most language models in our experiments have parameter sizes in the billions (e.g., 7B). Although we used LoRA fine-tuning to reduce the number of trainable parameters, inference still requires calculating all the parameters, resulting in longer inference times.
Appendix C Broader Impact
This paper presents work whose goal is to advance the field of LLMs. Our work evaluates the learning capabilities of large models through games and does not have negative societal impacts.